The aorta pumps blood away from the heart at about 40 cm/s and has a radius of about 1.0 cm. It then branches into many capillaries, each with a radius of about 5×10−4 cm carrying blood at a speed of 0.10 cm/s. How many capillaries are there? a- 1.6×109 b- 16 c 8.0×105 d- 1.6×105
Answer #1

we have to use equation of continuity

let N = number of capplilaries

(v*A)_1 = (v*A)_2
40cm/s * π(1.0cm)² = N[ 0.10cm/s * π*(5e-4cm)² ]
N = 1.6 x109
so option a is correct